# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 2.2 Boolean CSPs
  • Two-element semilattice: a two-element set together with the binary maximum operation with respect to one of the two possible orderings of the domain, e.g., $\mathbf{A}=(\{0,1\};\vee)$ where $\vee$ is the maximum operation. Here $\operatorname{Clo}(\mathbf{A})$ is the set of operations of the form $x_{i_{1}}\vee x_{i_{2}}\vee\dots\vee x_{i_{k}}$ and $\operatorname{Inv}(\mathbf{A})$ is the set of relations pp-definable from the singleton unary relations and the relation $x\vee y\vee\neg z$ (note that we abuse the notation and use $\vee$ both for the maximum operation and logical disjunction). Therefore, if $\mathbf{A}$ is a reduct of an algebra associated to a template $\mathbb{B}$, then the CSP over $\mathbb{B}$ can be reduced to Dual-Horn-3SAT (and thus to Horn-3SAT).
  • Two-element majority algebra: a two-element set together with the unique ternary operation that returns the majority of its arguments, e.g., $\mathbf{A}=(\{0,1\};\operatorname{maj})$, $\operatorname{maj}(a_{1},a_{2},a_{3})=1$ iff there are at least two $i$s with $a_{i}=1$. Here $\operatorname{Clo}(\mathbf{A})$ consists of all idempotent, monotone (compatible with the inequality relation $\leq$), and self-dual (i.e., compatible with the disequality relation $\neq$) operations and $\operatorname{Inv}(\mathbf{A})$ is the set of relations pp-definable from the binary relations $x\vee y$, $\neg x\vee y$, and $\neg x\vee\neg y$. Therefore, if $\mathbf{A}$ is a reduct of an algebra associated to a template $\mathbb{B}$, then the CSP over $\mathbb{B}$ can be reduced to 2SAT.
  • Affine Mal’cev algebra of a two-element group: a two-element set together with the ternary addition with respect to the unique group structure on the universe, e.g., $\mathbf{A}=(\{0,1\};x+y+z\pmod{2})$. Here $\operatorname{Clo}(\mathbf{A})$ is the set of operations of the form $x_{i_{1}}+x_{i_{2}}+\dots+x_{i_{k}}\pmod{2}$ with $k$ odd and $\operatorname{Inv}(\mathbf{A})$ is the set of relations pp-definable from the relations $x+y+z\equiv 1\pmod{2}$ and $x+y+z\equiv 0\pmod{2}$. Therefore, if $\mathbf{A}$ is a reduct of an algebra associated to a template $\mathbb{B}$, then the CSP over $\mathbb{B}$ can be solved by Gaussian elimination.
It follows from the full description of clones on a two-element set by Post [55] (and is not hard to verify) that every idempotent clone, except the clone consisting only of projections, contains one of the two semilattice operations or the majority operation or the ternary addition. Therefore, we have the following dichotomy for the CSP over $\mathbb{A}$ and an assigned algebra $\mathbf{A}$. Either
  • $\mathbf{A}$ contains only projections, in which case $\operatorname{Inv}(\mathbf{A})$ contains all relations, and then the CSP over $\mathbb{A}$ is NP-complete since any CSP, e.g. 3SAT, can be reduced to it; or
  • $\mathbf{A}$ contains some nonprojection, in which case one of three types of algebras above is a reduct of $\mathbf{A}$, and then the CSP over $\mathbb{A}$ is solvable in polynomial time.
@@ 2.3 Taylor algebras
For two-element idempotent structures, the necessary and sufficient condition for polynomial-time solvability is that $\mathbb{A}$ does not pp-define every finite structure or, in algebraic terms, $\mathbf{A}$ contains a nonprojection. For larger universes, this condition is no longer sufficient and pp-definability needs to be weakened to so called pp-interpretability. The algebraic counterpart goes as follows.
@@
Birkhoff’s HSP Theorem [13] connects the three main algebraic constructions (subalgebras, products, quotients) with identities, i.e., universally quantified equations. In particular, $\mathbf{A}$ is Taylor iff its operations satisfy some set of identities which are not satisfiable by projections. Several types of operations satisfying such identities generalize the three types in the previous subsections.
@@
  • A majority operation is a ternary operation $m$ satisfying $m(x,x,y)=m(x,y,x)=m(y,x,x)=x$ (for any $x,y$ in the universe). More generally, an $n$-ary near unanimity operation is an operation $f$ satisfying $f(x,x,\dots,x,y,x,x,\dots,x)=x$ for any position of $y$. An example for an odd $n$ is the the $n$-ary majority operation $\operatorname{maj}_{n}$ on $\{0,1\}$, that is, $\operatorname{maj}_{n}(a_{1},\dots,a_{n})=1$ iff the majority of the $a_{i}$ is 1.
@@
The two basic algorithmic ideas to efficiently solve a CSP are local propagation algorithms [33, 5] and finding a generating set of all solutions [12, 43].
We say that a CSP, or its assigned algebra $\mathbf{A}$, has bounded width if a local propagation algorithm correctly decides it. Examples include algebras with a semilattice or a majority operation, and a nonexample is an affine Mal’cev algebra of a nontrivial abelian group. The bounded width theorem [5] characterizes bounded width algebras as algebras $\mathbf{A}$ such that no quotient of a subalgebra of a power $\mathbf{A}$ is a nontrivial abelian algebra.
@@
An algebra $\mathbf{A}$ is abelian if the diagonal $\Delta_{A}=\{(a,a)\mathrel{:}a\in A\}$ is an equivalence class of a congruence of $\mathbf{A}^{2}$.
@@
The original proof of this result makes use of tame congruence theory, a well-developed theory of finite algebras that we have not mentioned yet. An alternative proof using absorption is available [8].
We now turn to the other algorithmic idea, finding a generating set of all solutions. It turns out [12] that $\mathbf{A}$ has “small” generating sets of powers (polynomial in the exponent) if and only if it has few subpowers, i.e., the number of subuniverses of $\mathbf{A}^{n}$ is $2^{O(n)}$. In such a case, the few subpowers algorithm [43] finds a generating set of solutions in polynomial time. Examples of algebras with few subpowers include algebras with a near unanimity or Mal’cev operation, and a nonexample is any nontrivial semilattice.
Algebras with few subpowers were characterized in [53] as those that do not have certain special subuniverses, called cube term blockers. It became clear that the concept is significant beyond this context [8, 61]. For this reason we prefer the terminology from the latter paper and call them projective subuniverses. In the definition, $\operatorname{Clo}_{n}(\mathbf{A})$ is the set of $n$-ary operations in $\operatorname{Clo}(\mathbf{A})$, and we use $a_{i}$ to denote the $i$-th component of a tuple ${\mathbf{a}}$.
@@
Let $\mathbf{A}$ be an algebra and $B\subseteq A$. We say that $B$ is a projective subuniverse if for every $f\in\operatorname{Clo}_{n}(\mathbf{A})$ there exists a coordinate $i$ of $f$ such that $f({\mathbf{a}})\in B$ whenever ${\mathbf{a}}\in A^{n}$ is such that $a_{i}\in B$.
Observe that a projective subuniverse of $\mathbf{A}$ is, indeed, a subuniverse.
Many of the algebraic concepts that we introduce (such as absorbing subuniverses or strongly projective subuniverses from Section 5) have a useful equivalent characterizations in terms of relations. Such a characterization for projective subuniverses is especially elegant and we state it here for comparison with characterizations of absorption in Subsection 5.2.
Proposition 2.6 (Lemma 3.2 in [53]).
Let $\mathbf{A}$ be an algebra and $B\subseteq A$. Then $B$ is a projective subuniverse of $\mathbf{A}$ if and only if, for every $n$, the relation $B(x_{1})\vee B(x_{2})\vee\dots\vee B(x_{n})$ is a subpower of $\mathbf{A}$.
@@ 3.1 Edges
We start by formally introducing the three types of edges used in Bulatov’s approach to the CSP. These three types are inspired by the three types of two-element algebras from Subsection 2.2. In the definition, $\operatorname{Sg}_{\mathbf{A}}(a,b)$ denotes the subalgebra of $\mathbf{A}$ generated by $\{a,b\}$.
@@
Let $\mathbf{A}$ be an algebra. A pair $(a,b)\in A^{2}$ is an edge if there exists a proper congruence $\theta$ on $\operatorname{Sg}_{\mathbf{A}}(a,b)$, called a witness for the edge, such that one of the following happens:
  • (semilattice edge) There is a term operation $f\in\operatorname{Clo}_{2}(\mathbf{A})$ acting as a join semilattice operation on $\{a/\theta,b/\theta\}$ with top element $b/\theta$, i.e. $f(a/\theta,b/\theta)$, $f(b/\theta,a/\theta)\subseteq b/\theta$.
  • (majority edge) There is a term operation $m\in\operatorname{Clo}_{3}(\mathbf{A})$ acting as a majority operation on $\{a/\theta,b/\theta\}$.
  • (abelian edge) The algebra $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ is abelian.
An edge $(a,b)$ is called minimal if for some maximal congruence $\theta$ witnessing the edge and every $a^{\prime},b^{\prime}\in A$ such that $(a,a^{\prime}),(b,b^{\prime})\in\theta$, we have $\operatorname{Sg}_{\mathbf{A}}(a^{\prime},b^{\prime})=\operatorname{Sg}_{\mathbf{A}}(a,b)$. Such $\theta$ is called a witness for this minimal edge.
A witnessing congruence $\theta$ for an edge $(a,b)$ necessarily separates $a$ and $b$, i.e., $(a,b)\not\in\theta$, since each congruence class of an idempotent algebra is a subuniverse. Moreover, if $\theta$ is a witness for an edge $(a,b)$, then any proper congruence of $\operatorname{Sg}_{\mathbf{A}}(a,b)$ containing $\theta$ witnesses the same edge.
@@
In order to make the concepts in this paper elegant and theorems more general, we deviate from the definition given in e.g. [22, 29]. There, majority edges have an additional requirement that the same congruence does not witness the semilattice type, and abelian edges (called affine) required the quotient to be an affine module. In Taylor algebras, abelian edges are the same as affine edges by Theorem 2.4. Outside Taylor algebras, it makes sense to separate abelian edges into two types: affine and sets whose only term operations are projections, as is done in [29].
@@
(The Connectivity Theorem [26, 22], link to proof) The directed graph formed by the edges of any algebra is (weakly) connected.
Another important technical result for the edge approach is that the operations appearing in Definition 3.1 can be significantly unified, see Theorem 7 in [26]. We give a much stronger unification result in Theorem 5.23.
@@ 3.2 Absorption
Let $\mathbf{A}$ be an algebra and $B\subseteq A$. We call $B$ an $n$-absorbing set of $\mathbf{A}$ if there is a term operation $t\in\operatorname{Clo}_{n}(\mathbf{A})$ such that $t({\mathbf{a}})\in B$ whenever ${\mathbf{a}}\in A^{n}$ and |$|\{i\mathrel{:}a_{i}\in B\}|\geq n-1$.
If, additionally, $B$ is a subuniverse of $\mathbf{A}$, we write $B\trianglelefteq_{n}\mathbf{A}$, or $B\trianglelefteq\mathbf{A}$ when the arity is not important.
We also say “$B$ absorbs $\mathbf{A}$ (by $t$)” in the situation of Definition 3.3. Of particular interest for us are $n$-absorbing subuniverses with $n=2$, e.g., $\{1\}$ in the semilattice ($(\{0,1\};\vee)$, or $n=3$, e.g., $\{0\}$ and $\{1\}$ in the two-element majority algebra on $\{0,1\}$.
The fundamental theorem for the absorption theory states that every proper subdirect linked subpower produces a nontrivial absorbing subuniverse. The required definitions are as follows. We say that $R$ is a subdirect product of $A_{1}$, …, $A_{n}$, and write $R\subseteq_{sd}A_{1}\times\dots\times A_{n}$, if $R\subseteq A_{1}\times\dots\times A_{n}$ and the projection of $R$ to any coordinate $i$ is full, i.e., equal to $A_{i}$. Similarly, we write $R\leq_{sd}\mathbf{A}_{1}\times\dots\times\mathbf{A}_{n}$ if $R$ is additionally a subuniverse of that product. If $R\subseteq_{sd}A\times C$, we call $R$ linked if it is connected when viewed as a bipartite graph between disjoint copies of $A$ and $C$.
Theorem 3.4 (The Absorption Theorem [4]).
Suppose $R\leq_{sd}\mathbf{A}\times\mathbf{C}$ is proper and linked, and $\mathbf{A}$ and $\mathbf{C}$ are Taylor. Then $\mathbf{A}$ or $\mathbf{C}$ has a nontrivial absorbing subuniverse.
We will often take advantage of one of the results of the absorption theory, the characterization of Taylor algebras by means of cyclic operations. An alternative proof is now available using Zhuk’s approach [61].
Theorem 3.5 ([4]).
@@
  • For every prime $p>|A|$, $\mathbf{A}$ has a term operation $t$ of arity $p$ which is cyclic.
@@ 3.3 Four types
The concept of a center is still evolving and it is not yet clear what the best version would be for general algebras. Our definition follows [60], although a more recent paper [61] made an adjustment motivated by this work. As we shall see in Theorem 5.10, the situation is much cleaner for minimal Taylor algebras.
@@
A subset $B\subseteq A$ is a center of $\mathbf{A}$ if there exists an algebra $\mathbf{C}$ (of the same signature) with no nontrivial 2-absorbing subuniverse and $R\leq_{sd}\mathbf{A}\times\mathbf{C}$ such that $B$ is the left center of $R$. The relation $R$ is called a witnessing relation. If $\mathbf{C}$ can be chosen Taylor, we call $B$ a Taylor center of $\mathbf{A}$.
An algebra $\mathbf{A}$ is polynomially complete if every operation on $A$ is in the clone generated by $\mathbf{A}$ together with the constant operations. This property can be equivalently phrased in terms of relations: $\mathbf{A}$ is polynomially complete if and only if it has no proper reflexive (that is, containing all the tuples $(a,a,\dots,a)$) irredundant (that is, no binary projection is the equality relation) subpowers.
@@
(The Four Types Theorem [60]) Let $\mathbf{A}$ be an algebra, then
@@
    $\mathbf{A}$ has a nontrivial 2-absorbing subuniverse, or
@@
    $\mathbf{A}$ has a nontrivial center (which is a Taylor center in the case where $\mathbf{A}$ is a Taylor algebra), or
@@
    $\mathbf{A}/\alpha$ is abelian for some proper congruence $\alpha$ of $\mathbf{A}$, or
@@
    $\mathbf{A}/\alpha$ is polynomially complete for some proper congruence $\alpha$ of $\mathbf{A}$.
We referred to four types of subuniverses whereas cases (c) and (d) talk about congruences – the subuniverses used in [60] are the equivalence classes of such congruences.
@@
Note, however, that Theorem 3.7 does not require that $\mathbf{A}$ is Taylor. If it is, then we get additional properties: centers are 3-absorbing by Proposition 4.3 and abelian algebras are term equivalent to affine modules by Theorem 2.4. For non-Taylor idempotent algebras, [61] suggests a similar five type theorem, which also follows immediately from our results.

### v6

@@ 2.2 Boolean CSPs
  • Two-element semilattice: a two-element set together with the binary maximum operation with respect to one of the two possible orderings of the domain, e.g., where $\vee$ is the maximum operation. Here $\operatorname{Clo}(\mathbf{A})$ is the set of operations of the form $x_{i_{1}}\vee x_{i_{2}}\vee\dots\vee x_{i_{k}}$ and $\operatorname{Inv}(\mathbf{A})$ is the set of relations pp-definable from the singleton unary relations and the relation $x\vee y\vee\neg z$ (note that we abuse the notation and use $\vee$ both for the maximum operation and logical disjunction). Therefore, if is a reduct of an algebra associated to a template $\mathbb{B}$, then the CSP over $\mathbb{B}$ can be reduced to Dual-Horn-3SAT (and thus to Horn-3SAT).
  • Two-element majority algebra: a two-element set together with the unique ternary operation that returns the majority of its arguments, e.g., , $\operatorname{maj}(a_{1},a_{2},a_{3})=1$ iff there are at least two $i$s with $a_{i}=1$. Here $\operatorname{Clo}(\mathbf{A})$ consists of all idempotent, monotone (compatible with the inequality relation $\leq$), and self-dual (i.e., compatible with the disequality relation $\neq$) operations and $\operatorname{Inv}(\mathbf{A})$ is the set of relations pp-definable from the binary relations $x\vee y$, $\neg x\vee y$, and $\neg x\vee\neg y$. Therefore, if is a reduct of an algebra associated to a template $\mathbb{B}$, then the CSP over $\mathbb{B}$ can be reduced to 2SAT.
  • Affine Mal’cev algebra of a two-element group: a two-element set together with the ternary addition with respect to the unique group structure on the universe, e.g., . Here $\operatorname{Clo}($ is the set of operations of the form $x_{i_{1}}+x_{i_{2}}+\dots+x_{i_{k}}\pmod{2}$ with $k$ odd and $\operatorname{Inv}(\mathbf{A})$ is the set of relations pp-definable from the relations $x+y+z\equiv 1\pmod{2}$ and $x+y+z\equiv 0\pmod{2}$. Therefore, if is a reduct of an algebra associated to a template $\mathbb{B}$, then the CSP over $\mathbb{B}$ can be solved by Gaussian elimination.
It follows from the full description of clones on a two-element set by Post [post41:clones] (and is not hard to verify) that every idempotent clone, except the clone consisting only of projections, contains one of the two semilattice operations or the majority operation or the ternary addition. Therefore, we have the following dichotomy for the CSP over $\mathbb{A}$ and an assigned algebra . Either
  • contains only projections, in which case $\operatorname{Inv}($ contains all relations, and then the CSP over $\mathbb{A}$ is NP-complete since any CSP, e.g. 3SAT, can be reduced to it; or
  • contains some nonprojection, in which case one of three types of algebras above is a reduct of , and then the CSP over $\mathbb{A}$ is solvable in polynomial time.
@@ 2.3 Taylor algebras
For two-element idempotent structures, the necessary and sufficient condition for polynomial-time solvability is that $\mathbb{A}$ does not pp-define every finite structure or, in algebraic terms, contains a nonprojection. For larger universes, this condition is no longer sufficient and pp-definability needs to be weakened to so called pp-interpretability. The algebraic counterpart goes as follows.
@@
Birkhoff’s HSP Theorem [Birkhoff] connects the three main algebraic constructions (subalgebras, products, quotients) with identities, i.e., universally quantified equations. In particular, is Taylor iff its operations satisfy some set of identities which are not satisfiable by projections. Several types of operations satisfying such identities generalize the three types in the previous subsections.
@@
  • A majority operation is a ternary operation $m$ which satisfies $m(x,x,y)=m(x,y,x)=m(y,x,x)=x$ (for any $x,y$ in the universe). More generally, an $n$-ary near unanimity operation is an operation $f$ satisfying $f(x,x,\dots,x,y,x,x,\dots,x)=x$ for any position of $y$. An example for an odd $n$ is the $n$-ary majority operation $\operatorname{maj}_{n}$ on $\{0,1\}$, that is, $\operatorname{maj}_{n}(a_{1},\dots,a_{n})=1$ iff the majority of the $a_{i}$ is 1.
@@
The two basic algorithmic ideas to efficiently solve a CSP are local propagation algorithms [Bulatov08:dualities, Barto14:local] and finding a generating set of all solutions [Berman10:varieties, Idziak10:few].
We say that a CSP, or its assigned algebra , has bounded width if a local propagation algorithm correctly decides it. Examples include algebras with a semilattice or a majority operation, and a nonexample is an affine Mal’cev algebra of a nontrivial abelian group. The bounded width theorem [Barto14:local] characterizes bounded width algebras as algebras such that no quotient of a subalgebra of a power $\mathbf{A}$ is a nontrivial abelian algebra.
@@
An algebra is abelian if the diagonal $\Delta_{A}=\{(a,a)\mathrel{:}a\in A\}$ is an equivalence class of a congruence of .
@@
The original proof of this result makes use of tame congruence theory, a well-developed theory of finite algebras that we have not mentioned yet. An alternative proof using absorption is available [Barto15:malcev].
We now turn to the other algorithmic idea, finding a generating set of all solutions. It turns out [Berman10:varieties] that has “small” generating sets of powers (polynomial in the exponent) if and only if it has few subpowers, i.e., the number of subuniverses of is $2^{O(n)}$. In such a case, the few subpowers algorithm [Idziak10:few] finds a generating set of solutions in polynomial time. Examples of algebras with few subpowers include algebras with a near unanimity or Mal’cev operation, and a nonexample is any nontrivial semilattice.
Algebras with few subpowers were characterized in [markovic12:cube-terms] as those that do not have certain special subuniverses, called cube term blockers. It became clear that the concept is significant beyond this context [Barto15:malcev, zhuk21:strong-subalgs]. For this reason we prefer the terminology from the latter paper and call them projective subuniverses. In the definition, $\operatorname{Clo}_{n}($ is the set of $n$-ary operations in $\operatorname{Clo}($, and we use $a_{i}$ to denote the $i$-th component of a tuple ${\mathbf{a}}$.
@@
Let be an algebra and $B\subseteq A$. We say that $B$ is a projective subuniverse if for every $f\in\operatorname{Clo}_{n}($ there exists a coordinate $i$ of $f$ such that $f({\mathbf{a}})\in B$ whenever ${\mathbf{a}}\in A^{n}$ is such that $a_{i}\in B$.
Observe that a projective subuniverse of is, indeed, a subuniverse.
Many of the algebraic concepts that we introduce (such as absorbing subuniverses or strongly projective subuniverses from Section 5) have useful equivalent characterizations in terms of relations. Such a characterization for projective subuniverses is especially elegant, and we state it here for comparison with characterizations of absorption in Subsection 5.2.
Proposition 2.6 (Lemma 3.2 in [markovic12:cube-terms]).
Let be an algebra and $B\subseteq A$. Then $B$ is a projective subuniverse of if and only if, for every $n$, the relation $B(x_{1})\vee B(x_{2})\vee\dots\vee B(x_{n})$ is a subpower of .
@@ 3.1 Edges
We start by formally introducing the three types of edges used in Bulatov’s approach to the CSP. These three types are inspired by the three types of two-element algebras from Subsection 2.2. In the definition, $\operatorname{Sg}(a,b)$ denotes the subalgebra of generated by $\{a,b\}$.
@@
Let be an algebra. A pair $(a,b)\in A^{2}$ is an edge if there exists a proper congruence $\theta$ on $\operatorname{Sg}(a,b)$, called a witness for the edge, such that one of the following happens:
  • (semilattice edge) There is a term operation $f\in\operatorname{Clo}_{2}($ acting as a join semilattice operation on $\{a/\theta,b/\theta\}$ with top element $b/\theta$, i.e. $f(a/\theta,b/\theta)$, $f(b/\theta,a/\theta)\subseteq b/\theta$.
  • (majority edge) There is a term operation $m\in\operatorname{Clo}_{3}($ acting as a majority operation on $\{a/\theta,b/\theta\}$.
  • (abelian edge) The algebra $\operatorname{Sg}(a,b)/\theta$ is abelian.
An edge $(a,b)$ is called minimal if for some maximal congruence $\theta$ witnessing the edge and every $a^{\prime},b^{\prime}\in A$ such that $(a,a^{\prime}),(b,b^{\prime})\in\theta$, we have $\operatorname{Sg}(a^{\prime},b^{\prime})=\operatorname{Sg}(a,b)$. Such $\theta$ is called a witness for this minimal edge.
A witnessing congruence $\theta$ for an edge $(a,b)$ necessarily separates $a$ and $b$, i.e., $(a,b)\not\in\theta$, since each congruence class of an idempotent algebra is a subuniverse. Moreover, if $\theta$ is a witness for an edge $(a,b)$, then any proper congruence of $\operatorname{Sg}(a,b)$ containing $\theta$ witnesses the same edge.
@@
In order to make the concepts in this paper elegant and theorems more general, we deviate from the definition given in e.g. [Bulatov04:graph, Bulatov20:local1]. There, majority edges have an additional requirement that the same congruence does not witness the semilattice type, and abelian edges (called affine) required the quotient to be an affine module. In Taylor algebras, abelian edges are the same as affine edges by Theorem 2.4. Outside Taylor algebras, it makes sense to separate abelian edges into two types: affine and sets whose only term operations are projections, as is done in [Bulatov20:local1].
@@
(The Connectivity Theorem [Bulatov16:graph, Bulatov04:graph], link to proof) The directed graph formed by the edges of any algebra is (weakly) connected.
Another important technical result for the edge approach is that the operations appearing in Definition 3.1 can be significantly unified, see Theorem 7 in [Bulatov16:graph]. We give a much stronger unification result in Theorem 5.23.
@@ 3.2 Absorption
Let be an algebra and $B\subseteq A$. We call $B$ an $n$-absorbing set of if there is a term operation $t\in\operatorname{Clo}_{n}($ such that $t({\mathbf{a}})\in B$ whenever ${\mathbf{a}}\in A^{n}$ and |$|\{i\mathrel{:}a_{i}\in B\}|\geq n-1$.
If, additionally, $B$ is a subuniverse of , we write $B\trianglelefteq_{n}$, or $B\trianglelefteq$ when the arity is not important.
We also say “$B$ absorbs (by $t$)” in the situation of Definition 3.3. Of particular interest for us are $n$-absorbing subuniverses with $n=2$, e.g., $\{1\}$ in the semilattice ($(\{0,1\};\vee)$, or $n=3$, e.g., $\{0\}$ and $\{1\}$ in the two-element majority algebra on $\{0,1\}$.
The fundamental theorem for the absorption theory states that every proper subdirect linked subpower produces a nontrivial absorbing subuniverse. The required definitions are as follows. We say that $R$ is a subdirect product of $A_{1}$, …, $A_{n}$, and write $R\subseteq_{sd}A_{1}\times\dots\times A_{n}$, if $R\subseteq A_{1}\times\dots\times A_{n}$ and the projection of $R$ to any coordinate $i$ is full, i.e., equal to $A_{i}$. Similarly, we write $R\leq_{sd}$ if $R$ is additionally a subuniverse of that product. If $R\subseteq_{sd}A\times C$, we call $R$ linked if it is connected when viewed as a bipartite graph between disjoint copies of $A$ and $C$.
Theorem 3.4 (The Absorption Theorem [Barto12:cyclic]).
Suppose $R\leq_{sd}$ is proper and linked, and and are Taylor. Then or has a nontrivial absorbing subuniverse.
We will often take advantage of one of the results of the absorption theory, the characterization of Taylor algebras by means of cyclic operations. An alternative proof is now available using Zhuk’s approach [zhuk21:strong-subalgs].
Theorem 3.5 ([Barto12:cyclic]).
@@
  • For every prime $p>|A|$, has a term operation $t$ of arity $p$ which is cyclic.
@@ 3.3 Four types
The concept of a center is still evolving, and it is not yet clear what the best version would be for general algebras. Our definition follows [Zhuk20:dichotomy], although a more recent paper [zhuk21:strong-subalgs] made an adjustment motivated by this work. As we shall see in Theorem 5.10, the situation is much cleaner for minimal Taylor algebras.
@@
A subset $B\subseteq A$ is a center of if there exists an algebra (of the same signature) with no nontrivial 2-absorbing subuniverse and $R\leq_{sd}$ such that $B$ is the left center of $R$. The relation $R$ is called a witnessing relation. If can be chosen Taylor, we call $B$ a Taylor center of .
An algebra is polynomially complete if every operation on $A$ is in the clone generated by together with the constant operations. This property can be equivalently phrased in terms of relations: is polynomially complete if and only if it has no proper reflexive (that is, containing all the tuples $(a,a,\dots,a)$) irredundant (that is, no binary projection is the equality relation) subpowers.
@@
(The Four Types Theorem [Zhuk20:dichotomy]) Let be an algebra, then
@@
    has a nontrivial 2-absorbing subuniverse, or
@@
    has a nontrivial center (which is a Taylor center in the case where is a Taylor algebra), or
@@
    is abelian for some proper congruence $\alpha$ of , or
@@
    is polynomially complete for some proper congruence $\alpha$ of .
We referred to four types of subuniverses whereas cases (c) and (d) talk about congruences – the subuniverses used in [Zhuk20:dichotomy] are the equivalence classes of such congruences.
@@
Note, however, that Theorem 3.7 does not require that is Taylor. If it is, then we get additional properties: centers are 3-absorbing by Proposition 4.3 and abelian algebras are term equivalent to affine modules by Theorem 2.4. For non-Taylor idempotent algebras, [zhuk21:strong-subalgs] suggests a similar five type theorem, which also follows immediately from our results.